# GeoPurify: A Data-Efficient Geometric Distillation Framework for Open-Vocabulary 3D Segmentation

### v1

@@ Abstract
Recent attempts to transfer features from 2D Vision–Language Models (VLMs) to 3D semantic segmentation expose a persistent trade-off. Directly projecting 2D features into 3D yields noisy and fragmented predictions, whereas enforcing geometric coherence necessitates costly training pipelines and large-scale, annotated 3D data. We argue that this limitation stems from the dominant segmentation-and-matching paradigm, which fails to reconcile 2D semantics with 3D geometric structure. The geometric cues are not eliminated during the 2D-to-3D transfer but remain latent within the noisy and view-aggregated features. To exploit this property, we propose GeoPurify that applies a small Student Affinity Network to purify 2D VLM-generated 3D point features using geometric priors distilled from a 3D self-supervised teacher model. During inference, we devise a Geometry-Guided Pooling module to further denoise the point cloud and ensure the semantic and structural consistency. Benefiting from latent geometric information and the learned affinity network, GeoPurify effectively mitigates the trade-off and achieves superior data efficiency. Extensive experiments on major 3D benchmarks demonstrate that GeoPurify achieves or surpasses state-of-the-art performance while utilizing only $\sim$1.5% of the training data. Our codes and checkpoints are available in https://github.com/tj12323/GeoPurify.
footnotetext: $\dagger$ Corresponding Authors: yi.bin@hotmail.com
@@
In the 2D-to-3D transfer, the 2D models typically follow a “Segmentation and Matching” paradigm (Li et al., 2022; Ghiasi et al., 2022; Wang et al., 2024a), a brittle two-step process that first segments objects and then classifies them, treating geometry and semantics as disconnected problems. This sharply contrasts with biological perception, where recognizing the form of an object and understanding its meaning are intrinsically linked, following a “Segmentation as Understanding” paradigm. We argue that machine perception should emulate this approach, where meaningful segmentation emerges as a property of deeper contextual understanding rather than as a precursor to recognition. However, as shown in Figure 1-(b), 2D VLM features ($F_{sem}$) are semantically rich but geometrically inconsistent, resulting fragments and shape distortion, whereas priors from 3D self-supervised models ($G_{geo}$) are geometrically reliable but lack open-vocabulary semantics, as illustrated in Figure 1-(c). Inspired by Wysoczańska et al. (2024), we hypothesize that transferring 2D features into 3D does not destroy the geometric information, but rather renders it latent, suggesting that we may recover latent structure more efficiently instead of learning 3D geometry from scratch.
@@
Motivated by this hypothesis, we present GeoPurify, a “Segmentation as Understanding” framework designed to recover latent geometric structure and produce robust 3D representations. We first leverage a one-for-all VLM to generate semantic-rich and geometric-aware 3D features from multiple 2D views. To address the noise and fragmentation inherent in VLM-generated point clouds, we propose Geometric Contrastive Distillation, where a student module learns latent geometric affinities from noisy features under the guidance of a 3D teacher’s structural priors. By exploiting the pre-existing geometric knowledge, GeoPurify eliminates the need for large-scale manual annotation and can be trained using only a small set of unlabeled 3D scans. At inference, a Geometry-Guided Pooling module is designed to denoise and consolidate the features, producing unified 3D representations that are both semantically rich and geometrically coherent. Experiments on multiple popular 3D semantic segmentation datasets demonstrate that GeoPurify achieves comparable or superior performance to state-of-the-art methods while using only $\sim$1.5% of the training data, effectively resolving the trade-off that limits prior approaches. In summary, our contributions are:
  • We reframe open-vocabulary 3D segmentation from “Segmentation and Matching” to “Segmentation as Understanding”. By leveraging a generalist VLM, we establish a richer semantic foundation in a unified visual-linguistic space, overcoming the low semantic ceiling of prior “localize-then-recognize” approaches.
  • We introduce GeoPurify, a data-efficient framework built on a novel Geometric Contrastive Distillation mechanism. Its core components are: a Student Affinity Network to learn latent geometric affinities from unlabeled 3D scans, and Geometry-Guided Pooling that uses these affinities to denoise features and ensure structural consistency at inference.
@@ 2 Related Work
To overcome closed-set limitations, open-vocabulary 3D scene understanding methods transfer rich semantics from 2D Vision-Language Models (VLMs). The core challenge is rectifying geometric inconsistencies that arise when projecting multi-view 2D features into a unified 3D space. Current solutions often tackle this with extensive supervision. Some perform simple training-free feature aggregation, but this yields noisy and fragmented 3D features (Jatavallabhula et al., 2023; Lu et al., 2023; Peng et al., 2023). While subsequent post-processing methods like PoVo (Mei et al., 2025) attempt to refine these initial features, they often fail to enforce consistency for complex or occluded objects. The dominant approach is large-scale knowledge distillation, where a 3D backbone is trained on noisy pseudo-labels from 2D projections, requiring massive datasets to implicitly learn geometric priors (Chen et al., 2023b; Peng et al., 2023; Li et al., 2024; Wang et al., 2024c; Li et al., 2025; Wang et al., 2024). Others attempt direct 3D-text alignment, but typically rely on coarse, region-level text supervision that lacks the granularity for dense prediction tasks (Ding et al., 2023; Yang et al., 2024; Wang et al., 2025; Jiang et al., 2024). Ultimately, while techniques like AFOV (Sun et al., 2025) enhance annotation efficiency, they still share a fundamental reliance on massive datasets to implicitly resolve geometric ambiguities through improved supervision. This reveals a critical gap for a data-efficient method that explicitly enforces geometric coherence during the 2D-to-3D semantic transfer.
@@ 3.2 Semantic Initialization from a Generalist VLM
To obtain 3D representations enriched with semantic priors, we project RGB inputs into the 3D point space using a frozen 2D VLM and merge multiple views to capture structural cues. Prior methods often follow a Segmentation and Matching paradigm, as in LSeg (Li et al., 2022), OpenSeg (Ghiasi et al., 2022), and SAM+CLIP frameworks (Wang et al., 2024a), which adopt a Localize then Recognize pipeline. They first group pixels into regions using segmentation backbones and then match these clusters to predefined text labels. While effective in closed-set settings, this approach produces highly discriminative yet task-specific features, limiting their semantic ceiling and reducing their capacity to capture the generative and associative richness needed in open-world 3D environments. To overcome this limitation, we adopt X-Decoder (Zou et al., 2023) as our frozen 2D VLM backbone ($\Psi_{2D}$). Instead of isolating segmentation as a downstream task, X-Decoder constructs a unified visual-linguistic embedding space that supports diverse tasks such as segmentation, retrieval, and dense captioning. Its dual-query architecture, which learns from both latent visual queries and explicit text queries, makes segmentation an emergent property of deeper contextual understanding. The resulting feature space is inherently generative and associative, optimized to answer the holistic question: What is depicted, where is it located, and how can it be described in language? These richer representations provide a higher semantic ceiling and yield stronger 2D priors for 3D scene understanding.
@@ 4.2 Quantitative Results
As shown in Table 1, GeoPurify demonstrates exceptional performance by leveraging a curated data subset of only $\sim$1.5%. To rigorously validate that our method’s efficacy stems from its architecture rather than just the data selection, we retrained a leading competitor, CUA-O3D (Li et al., 2025), on the exact same compact subset. Under these identical data-efficient conditions, CUA-O3D’s performance drops sharply to an 18.1 mIoU on ScanNetV2. In stark contrast, GeoPurify achieves an mIoU of 55.1. This significant performance gap underscores our method’s unique ability to distill robust priors from a small but semantically dense data source.
Remarkably, our model not only excels in this low-data regime but also rivals methods trained on the full dataset. The advantage is even more pronounced in the mean Accuracy (mAcc) metric, where GeoPurify attains 72.5 on ScanNetV2 and 62.4 on Matterport3D, substantially outperforming fully-trained competitors. This divergence between high mAcc and competitive mIoU is an intentional outcome of our design. The superior mAcc is driven by Geometry-Guided Pooling, which propagates VLM-initialized core labels across geometrically coherent regions, enforcing intra-object consistency and boosting recall. This mechanism, however, introduces a deliberate trade-off: by prioritizing semantic purity within objects, it can cause minor “semantic bleeding” at boundaries, slightly degrading the boundary precision to which mIoU is sensitive. For a detailed qualitative analysis, including visual examples, please refer to Appendix C.
@@
Type Method
3D Net
Training Data
ScanNetV2 Matterport3D
@@ 4.3 Ablation Studies
Geometric Purification
2D Semantic Backbone
Hybrid Contrastive Sampling
Pooling Iterations ($T$)
Table 4: Ablation studies of GeoPurify on the ScanNetV2 validation set. We analyze the impact of our core geometric purification module, the choice of 2D backbone, the contrastive sampling strategy, and the number of pooling iterations ($T$). Our default configuration is highlighted in bold.
@@ B.2 Cross-Dataset Generalization
Appendix C Extended Qualitative Analysis
@@
C.1 Visualization of Intermediate Mechanisms
@@ C.2 Qualitative Results
C.3 Failure Case Analysis
@@
Appendix D The Use of Large Language Models

### v2

@@ Abstract
Recent attempts to transfer features from 2D Vision–Language Models (VLMs) to 3D semantic segmentation expose a persistent trade-off. Directly projecting 2D features into 3D yields noisy and fragmented predictions, whereas enforcing geometric coherence necessitates costly training pipelines and large-scale, annotated 3D data. We argue that this limitation stems from the dominant segmentation-and-matching paradigm, which fails to reconcile 2D semantics with 3D geometric structure. The geometric cues are not eliminated during the 2D-to-3D transfer but remain latent within the noisy and view-aggregated features. To exploit this property, we propose GeoPurify that applies a small Student Affinity Network to purify 2D VLM-generated 3D point features using geometric priors distilled from a 3D self-supervised teacher model. During inference, we devise a Geometry-Guided Pooling module to further denoise the point cloud and ensure the semantic and structural consistency. Benefiting from latent geometric information and the learned affinity network, GeoPurify effectively mitigates the trade-off and achieves superior data efficiency. Extensive experiments on major 3D benchmarks demonstrate that GeoPurify achieves or surpasses state-of-the-art performance while utilizing only $\sim$1.5% of the training data.
      Code Repository: https://github.com/tj12323/GeoPurify
@@
In the 2D-to-3D transfer, the 2D models typically follow a “Segmentation and Matching” paradigm (Li et al., 2022; Ghiasi et al., 2022; Wang et al., 2024a), a brittle two-step process that disconnects geometry and semantics. We argue that machine perception should instead follow a “Segmentation as Understanding” paradigm, where recognizing an object’s form and understanding its meaning are intrinsically linked. However, simply adopting a generalist VLM that follows this paradigm is insufficient. As shown in Figure 1-(b), 2D VLM features ($F_{sem}$) are semantically rich but geometrically inconsistent, resulting fragments and shape distortion, whereas priors from 3D self-supervised models ($G_{geo}$) are geometrically reliable but lack open-vocabulary semantics, as illustrated in Figure 1-(c). Inspired by Wysoczańska et al. (2024), we hypothesize that transferring 2D features into 3D does not destroy the geometric information, but rather renders it latent, suggesting that we may recover latent structure more efficiently instead of learning 3D geometry from scratch.
@@
Motivated by this hypothesis, we present GeoPurify, a data-efficient framework designed to recover latent geometric structure from noisy semantic features and produce robust 3D representations. We first leverage a one-for-all VLM to generate semantic-rich and geometric-aware 3D features from multiple 2D views. To address the noise and fragmentation inherent in VLM-generated point clouds, we propose Geometric Contrastive Distillation, where a student module learns latent geometric affinities from noisy features under the guidance of a 3D teacher’s structural priors. By exploiting the pre-existing geometric knowledge, GeoPurify eliminates the need for large-scale manual annotation and can be trained using only a small set of unlabeled 3D scans. At inference, a Geometry-Guided Pooling module is designed to denoise and consolidate the features, producing unified 3D representations that are both semantically rich and geometrically coherent. Experiments on multiple popular 3D semantic segmentation datasets demonstrate that GeoPurify achieves comparable or superior performance to state-of-the-art methods while using only $\sim$1.5% of the training data, effectively resolving the trade-off that limits prior approaches.
In summary, our contributions are:
  • We introduce GeoPurify, a data-efficient framework built on the hypothesis that beyond their semantic richness, VLM-projected features also embed a latent 3D geometric structure. GeoPurify demonstrates that recovering this latent structure is a more data-efficient path to open-vocabulary 3D segmentation.
  • Our framework first leverages a generalist VLM to establish a rich semantic foundation via “Segmentation as Understanding”, overcoming the low semantic ceiling of prior “Segmentation and Matching” approaches . It then introduces a Geometric Contrastive Distillation mechanism to learn latent geometric affinities from unlabeled 3D scans and a Geometry-Guided Pooling module that uses these affinities to denoise features and ensure structural consistency at inference.
@@ 2 Related Work
To overcome closed-set limitations, open-vocabulary 3D scene understanding methods transfer rich semantics from 2D Vision-Language Models (VLMs). The core challenge is rectifying geometric inconsistencies that arise when projecting multi-view 2D features into a unified 3D space. Current solutions often tackle this with extensive supervision. Some perform simple training-free feature aggregation, but this yields noisy and fragmented 3D features (Jatavallabhula et al., 2023; Lu et al., 2023; Peng et al., 2023; Nguyen et al., 2024). While subsequent post-processing methods like PoVo (Mei et al., 2025) attempt to refine these initial features, they often fail to enforce consistency for complex or occluded objects. The dominant approach is large-scale knowledge distillation, where a 3D backbone is trained on noisy pseudo-labels from 2D projections, requiring massive datasets to implicitly learn geometric priors (Chen et al., 2023b; Peng et al., 2023; Li et al., 2024; Wang et al., 2024c; Li et al., 2025; Wang et al., 2024; Xu et al., 2024). Others attempt direct 3D-text alignment, but typically rely on coarse, region-level text supervision that lacks the granularity for dense prediction tasks (Ding et al., 2023; Yang et al., 2024; Wang et al., 2025; Jiang et al., 2024). Recent works such as OV3D (Jiang et al., 2024) and PGOV3D (Zhang et al., 2025) advance VLM-based understanding pipelines but typically rely on large-scale data for alignment or curriculum learning, missing the efficiency of explicitly recovering latent geometric affinities. Ultimately, while techniques like AFOV (Sun et al., 2025) enhance annotation efficiency, they still share a fundamental reliance on massive datasets to implicitly resolve geometric ambiguities through improved supervision. This reveals a critical gap for a data-efficient method that explicitly enforces geometric coherence during the 2D-to-3D semantic transfer.
@@ 3.2 Semantic Initialization from a Generalist VLM
To obtain 3D representations enriched with semantic priors, we project RGB inputs into the 3D point space (constructed by aggregating multi-view projections, without necessitating external raw point clouds at inference) using a frozen 2D VLM and merge multiple views to capture structural cues. Prior methods often follow a Segmentation and Matching paradigm, as in LSeg (Li et al., 2022), OpenSeg (Ghiasi et al., 2022), and SAM+CLIP frameworks (Wang et al., 2024a), which adopt a Localize then Recognize pipeline. They first group pixels into regions using segmentation backbones and then match these clusters to predefined text labels. While effective in closed-set settings, this approach produces highly discriminative yet task-specific features, limiting their semantic ceiling and reducing their capacity to capture the generative and associative richness needed in open-world 3D environments. To overcome this limitation, we adopt X-Decoder (Zou et al., 2023) as our frozen 2D VLM backbone ($\Psi_{2D}$). Instead of isolating segmentation as a downstream task, X-Decoder constructs a unified visual-linguistic embedding space that supports diverse tasks such as segmentation, retrieval, and dense captioning. Its dual-query architecture, which learns from both latent visual queries and explicit text queries, makes segmentation an emergent property of deeper contextual understanding. The resulting feature space is inherently generative and associative, optimized to answer the holistic question: What is depicted, where is it located, and how can it be described in language? These richer representations provide a higher semantic ceiling and yield stronger 2D priors for 3D scene understanding.
@@ 4.2 Quantitative Results
As shown in Table 1, GeoPurify demonstrates exceptional performance by leveraging a curated data subset of only $\sim$1.5%. To rigorously validate that our method’s efficacy stems from its architecture rather than just the data selection, we retrained a leading competitor, CUA-O3D (Li et al., 2025), on the exact same compact subset. Under these identical data-efficient conditions, CUA-O3D’s performance drops sharply to an 18.1 mIoU on ScanNetV2. In stark contrast, GeoPurify achieves an mIoU of 55.1. We attribute this robustness to our decoupled design. Standard methods typically attempt to learn entangled geo-semantic representations from scratch, failing to generalize without sufficient data. In contrast, GeoPurify relies on the frozen VLM for semantics and trains the student network solely to recover latent geometric structure. Learning these geometric grouping rules requires significantly less data, allowing our method to function as an efficient “geometric denoiser” that maintains high performance.
Remarkably, our model not only excels in this low-data regime but also rivals methods trained on the full dataset. The advantage is even more pronounced in the mean Accuracy (mAcc) metric, where GeoPurify attains 72.5 on ScanNetV2 and 62.4 on Matterport3D, substantially outperforming fully-trained competitors. This divergence between high mAcc and competitive mIoU is an intentional outcome of our design. The superior mAcc is driven by Geometry-Guided Pooling, which propagates VLM-initialized core labels across geometrically coherent regions, enforcing intra-object consistency and boosting recall. This mechanism, however, introduces a deliberate trade-off: by prioritizing semantic purity within objects, it can cause minor “semantic bleeding” at boundaries, slightly degrading the boundary precision to which mIoU is sensitive. For a detailed qualitative analysis, including visual examples, please refer to Appendix D.
@@
Type Method 3D Net Training Data ScanNetV2 Matterport3D
@@ 4.3 Ablation Studies
Table 4: Ablation studies of GeoPurify on the ScanNetV2 validation set. We analyze the impact of our core geometric purification module, the choice of 2D backbone, the contrastive sampling strategy, the number of pooling iterations ($T$), and the training subset size. Our default configuration is highlighted in bold.
@@
Training Subset Size 10 scenes 54.7 72.4
20 scenes (Ours) 55.1 72.5
30 scenes 55.1 72.5
50 scenes 55.0 72.5
Geometric Purification
2D Semantic Backbone
Hybrid Contrastive Sampling
Pooling Iterations ($T$)
Training Subset Size
We evaluate the method’s sensitivity to data quantity by training on subsets of 10, 20, 30, and 50 scenes. As shown in Table 4, performance improves markedly from 10 to 20 scenes but plateaus thereafter. This trend validates our design principle: unlike methods learning complex semantic mappings, our student network distills class-agnostic geometric affinities (e.g., surface continuity) which are repetitive and universal. Consequently, a small subset (20 scenes) provides sufficient structural diversity for convergence. The observed plateau indicates that our framework successfully recovers the latent structure available to refine the current VLM features, suggesting performance is primarily bounded by the VLM’s semantic quality rather than the quantity of geometric training data.
@@ 5 Conclusion
Acknowledgments
This work is supported by the Central Guidance on Local Science and Technology Development Fund of Shanghai City (YDZX20253100002004), National Key Research and Development Program of China (2025YFF0522500), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM116), and the Fundamental Research Funds for the Central Universities.
@@ B.2 Cross-Dataset Generalization
Appendix C Subset Selection Strategy
To ensure that the efficacy of GeoPurify is not an artifact of the specific data sampling method, we evaluate its robustness against two fully annotation-free alternatives: random uniform selection and unsupervised clustering based on spatial and colorimetric features. As reported in Table 11, the model trained on randomly selected scenes achieves an mIoU of 54.6%, representing a negligible deviation from the heuristic baseline. Similarly, the unsupervised clustering strategy yields 54.9% mIoU. These findings confirm that the observed performance gains are intrinsic to the proposed geometric purification mechanism and do not rely on specific high-quality samples or global statistical guidance.
Table 11: Robustness analysis of subset selection strategies on ScanNetV2 ($\sim$1.5% data). The comparison demonstrates that GeoPurify maintains consistent performance across varying selection methodologies.
Selection Strategy mIoU mAcc
Heuristic (Ours) 55.1 72.5
Random 54.6 72.2
Coordinate and Color Clustering 54.9 72.3
Appendix D Extended Qualitative Analysis
@@
D.1 Visualization of Intermediate Mechanisms
@@ D.2 Qualitative Results
D.3 Failure Case Analysis
@@
Appendix E The Use of Large Language Models